OpenStudy (anonymous):

(1+tanθ)/(sinθ+cosθ)=secθ

OpenStudy (anonymous):

what to do?

OpenStudy (anonymous):

\[\frac{1+\tan \theta}{\sin \theta+\cos \theta}=\sec \theta\]

OpenStudy (anonymous):

verify indentity

OpenStudy (anonymous):

(1+tanθ)/(sinθ+cosθ) = 1/cosθ cosθ/(sinθ +cosθ) = 1/(1+tanθ) tanθ = sinθ/cosθ 1 + tanθ = 1 + sinθ/cosθ = (cosθ + sinθ)/cosθ 1/(1+tanθ) = cosθ/(cosθ + sinθ) which is what is in the left side :)

OpenStudy (anonymous):

put any value of theta and you will se that the rhs and lhs are same

OpenStudy (anonymous):

RHS\[\sec \theta=\frac{\sin \theta + \cos \theta}{\sin \theta +\cos \theta}\sec \theta\]\[=\frac{1+\tan \theta}{\sin \theta + \cos \theta}=RHS\]

OpenStudy (anonymous):

you could actually do it without my first steps .. : 1 + tanθ = 1 + sinθ/cosθ = (cosθ + sinθ)/cosθ so put it in the original : ( (cosθ + sinθ)/cosθ) / ( sinθ + cosθ) =1 /cosθ ( (cosθ + sinθ)/(cosθ( sinθ + cosθ))) = 1/cosθ then 1/cosθ = 1/cosθ whatever you prefer

OpenStudy (anonymous):

i think that u got it..

OpenStudy (anonymous):

keep right side, solve left side?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

you can do that also

OpenStudy (anonymous):

\[\frac{1+\tan \theta}{\sin \theta+\cos \theta }=\frac{\cos \theta (1 +\tan \theta)}{\cos \theta (\sin \theta +\cos \theta)}\]\[\frac{\sin \theta + \cos \theta}{\cos \theta(\sin \theta +\cos \theta)}=1/\cos \theta =\sec \theta =RHS\]

OpenStudy (anonymous):

I don't get

OpenStudy (anonymous):

\[\frac{\cos \theta+\sin \theta}{\frac{\cos \theta}{\sin \theta+\cos \theta}}\]I get from here

OpenStudy (anonymous):

\[=\frac{(\cos \theta+\sin \theta)^2}{\cos \theta}\]

OpenStudy (anonymous):

what are you doing????

OpenStudy (anonymous):

start from any one of the sides..either RHS or LHS

OpenStudy (anonymous):

I do 1+tan

OpenStudy (anonymous):

take 1+tan/(sin + cos)

OpenStudy (anonymous):

multiply num and den with cos and see my previous post

OpenStudy (anonymous):

now got it?